OpenStudy (anonymous):

if side AD= 18.02 cm and side AB=20 cm, calculate the length of side AC. A. 16.4 cm B. 15.5 cm C. 20 cm D. 22.2 cm hangon a sec I gota draw it:)

OpenStudy (anonymous):



OpenStudy (anonymous):

@jennychan12

OpenStudy (jennychan12):

ohmigosh. i'm sorry. i suck at this, and i hated geometry. so sorry, no can do.

OpenStudy (anonymous):

do u know someone who can?

OpenStudy (jennychan12):

sorry, no. :(

OpenStudy (anonymous):

@Luis_Rivera please please help!!?

OpenStudy (jennychan12):

maybe this will help? http://www.mathwarehouse.com/geometry/circle/

OpenStudy (anonymous):

@phi

OpenStudy (jennychan12):

ok so i THINK it's AB^2 = AD(CD+AD) judging by the link i gave u.

OpenStudy (jennychan12):

did it help? @katlin95 ? did u get ur answer?: if so, good :)

OpenStudy (anonymous):

I dont know the number to put it for AB^2=AD(CD+AD) what did u get?

OpenStudy (phi):

you said I dont know the number to put it for AB^2=AD(CD+AD) side AD= 18.02 cm and side AB=20 cm, you should learn the idea of replacing a letter (or name) with its number. they tell you AB=20. everywhere you see AB, replace it with 20 in the formula also, replace AD with 18.02 in AB^2=AD(CD+AD) it really is worth understanding this idea: replace letters with numbers... 20^2 = 18.02*(CD + 18.02) I would now solve for CD (CD is a length you are looking for, so you can add it to AD and so get AC) if we call CD x you get this equation 20*20= 18.02*(x + 18.02) solve for x: multiply 18.02 times each term in side the parens. as the first step....